OpenStudy (anonymous):

if m<1=33% and m<2=18%,what is m<3?

OpenStudy (anonymous):

100%-33%-18%=51%

OpenStudy (anonymous):

15 is not a option for a answer in my homework,sorry it only one circle,the circle symbol of degress

OpenStudy (anonymous):

i mean degrees

OpenStudy (anonymous):

a:129 b:114 c:162 d:147 all of the option have the circle symbol of degrees

OpenStudy (anonymous):

https://app37.studyisland.com/pics/120549pic10A.png

OpenStudy (p0sitr0n):

how about 180-18-33?

OpenStudy (anonymous):

i dont thats a option for a answer

OpenStudy (ranga):

It is a triangle. Three angles of a triangle add to 180 degrees. The sum of the given two angles is 18 + 33 = 51 So the third angle must be: 180 - 51 = ?

OpenStudy (anonymous):

149 i guess?

OpenStudy (ranga):

you just have to subtract 51 from 180.

OpenStudy (anonymous):

thats what i did its 149 am i right?

OpenStudy (ranga):

really? you are subtracting 51 from 200. subtract 51 from 180.

OpenStudy (anonymous):

129?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

okay thanks.

OpenStudy (ranga):

yw.

OpenStudy (anonymous):

I have another question

OpenStudy (ranga):

post it.

OpenStudy (anonymous):

which shows the expression below in simpliest form? (3.2 x 10^6) divided by (4x10^3)

OpenStudy (anonymous):

the options for the answer are: a:-0.8 x 10^2 b:8x10^3 c:8x10^4 d:8x10^2

OpenStudy (ranga):

\[\frac{ 3.2 \times 10^6 }{ 4 \times 10^3 } = \frac{ 3.2 }{ 4 } \times \frac{ 10^6 }{ 10^2 }\] Simplify.

OpenStudy (ranga):

brb. I'm answering another question.

OpenStudy (anonymous):

okay

OpenStudy (ranga):

If you divide 3.2/4 you get 0.8 10^6/10^2 = 10^(6-2) = 10^4 So 0.8 x 10^4. But in scientific notation the first number should be between 1 and 10. So, 0.8 x 10^4. = 0.8 x 10 x 10^3 = 8 x 10^3